OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. 5. (1 pt) Which answer describes the solution to each equation? Column A Column B 1. 5x + 7 = 1/2 (10x +12) 2. 1/9 (18x + 27) = 1+ x -4 - 3x 3. x + 3/4 = 3/4x + 3/4 4. 4 + 3x = 2x - 10 A. one solution B. no solution C. an infinite number of solutions

OpenStudy (anonymous):



OpenStudy (mathmale):

Please, first do all you can yourself to work through the problem and to narrow down your choice of answers. What have you done so far? What are your thoughts about these problems?

OpenStudy (anonymous):

ive gotton a b and c

OpenStudy (texaschic101):

1. 5x + 7 = 1/2(10x + 12) 5x + 7 = 5x + 6 false.....no solutions 2. 1/9(18x + 27) = 1 + x - 4 + 3x 2x + 3 = -3 + 4x 2x - 4x = -3 - 3 -2x = -6 x = 3 one solution 3. x + 3/4 = 3/4x + 3/4 x - 3/4x = 3/4 - 3/4 4/4x - 3/4x = 0 1/4x = 0 x = 0 one solution 4. 4 + 3x = 2x - 10 3x - 2x = -10 - 4 x = -14 one solution